OpenStudy (anonymous):

If the population of bacteria doubles in 4 hours how long will it take for it to triple its original size

OpenStudy (anonymous):

\[2^{\frac{t}{4}}=3\] \[\frac{t}{4}=\frac{\ln(3)}{\ln(2)}\] \[t=\frac{4\ln(3)}{\ln(2)}\]

OpenStudy (anonymous):

6.34 hours?

OpenStudy (anonymous):

i don't know, need a calculator

OpenStudy (anonymous):

yes, that is what i get

OpenStudy (anonymous):

k thanks!

OpenStudy (anonymous):

can you help with this one too?

OpenStudy (anonymous):

if a sales of a company increases exponentially by 50% from 1985 to 1990, by what percent will the sales increase from 1990 to 1996